# DiffECG: A Versatile Probabilistic Diffusion Model for ECG Signals Synthesis

### v2

@@ 5 Conclusion
In this paper, we presented the first versatile conditional diffusion framework for ECG synthesis that can perform three different tasks: heartbeats generation, imputation and forecasting. The obtained results demonstrated the effectiveness of our approach, as well as its ability to enhance state-of-the-art classifiers’ performance. For future work, we plan to investigate the combination of diffusion models with adversarial training to further enhance ECG synthesis. Additionally, we aim to extend our approach to generate others classes of arrhythmia and multi-lead ECG signals, while also considering the synthesis of other physiological signals.

### v3

@@ V Conclusion
In this paper, we presented the first versatile conditional diffusion framework for ECG synthesis that can perform three different tasks: heartbeats generation, imputation, and forecasting. The obtained results demonstrated the effectiveness of our approach, as well as its ability to enhance state-of-the-art classifiers’ performance. For future work, we plan to investigate the combination of diffusion models with adversarial training to further enhance ECG synthesis. Additionally, we aim to extend our approach to generate other classes of arrhythmia and multi-lead ECG signals, while also considering the synthesis of other physiological signals.